OpenStudy (anonymous):

Help Please!! Use Laplace transforms to solve the initial value problem x''+4x'+8x=2e^(-t) ; x(0)=0 and x'(0)=4

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

this is what I got so far

OpenStudy (anonymous):

after this I am stuck :(

OpenStudy (anonymous):

and this is the table I used: http://tutorial.math.lamar.edu/pdf/Laplace_Table.pdf

OpenStudy (turingtest):

use partial fractions on the \[\frac2{s(s+1)(s+4)}\]part and I think it should be pretty straightforward to do inverse Laplace

OpenStudy (anonymous):

but what I do about the 4/s^2+4s

OpenStudy (turingtest):

factor out the s and that is what I wrote

OpenStudy (turingtest):

oh I see; same thing though factor out s and do partial fractions

OpenStudy (anonymous):

but then you get 4/(s(s+4)-->4/s[1/s+4)--> 4e^(-4t)

OpenStudy (anonymous):

is that right?

OpenStudy (turingtest):

\[\frac4{s^2+4s}=\frac As+\frac B{s+4}\implies A(s+4)+Bs=4\]\[A=1~~~,B=-1\]

OpenStudy (anonymous):

oh so you do partial fractions with both of them

OpenStudy (turingtest):

that's what I'd do, yeah

OpenStudy (anonymous):

and for the other one it's A(s^2+4s) B(s+1) right

OpenStudy (turingtest):

\[\mathcal L^{-1}\{\frac1s\}=1\]\[-\mathcal L^{-1}\{\frac1{s+4}\}=-e^{-4t}\]

OpenStudy (turingtest):

for the other one it would be...

OpenStudy (turingtest):

\[\frac2{s(s+1)(s+4)}=\frac As+\frac B{s+1}+\frac C{s+4}\]

OpenStudy (turingtest):

\[A(s+1)(s+4)+Bs(s+4)+Cs(s+1)=2\]

OpenStudy (anonymous):

then you get As^2+5As+4A+Bs^2+4Bs+Cs^2+C=2

OpenStudy (anonymous):

right

OpenStudy (turingtest):

not necessary I don't think plug in s=-1,-4,and 0 and I think you can get the values quickly

OpenStudy (anonymous):

so you dont need ti distribute it?

OpenStudy (turingtest):

not necessarily how did I get the other answers? I'll show you...

OpenStudy (anonymous):

i didnt get the answers yet i thiught you had to distribute it then ge tthe answers but now i see that its unnecessary

OpenStudy (turingtest):

\[\frac4{s^2+4s}=\frac As+\frac B{s+4}\implies A(s+4)+Bs=4\]now plug in\[s=0:A(0+4)+B(0)=4 \implies 4A=4\implies A=1\]

OpenStudy (anonymous):

B=-2/3 C=1/6

OpenStudy (anonymous):

and A=1/2

OpenStudy (anonymous):

thats what I got

OpenStudy (turingtest):

this is the second term:\[\frac4{s^2+4s}=\frac As+\frac B{s+4}\implies A(s+4)+Bs=4\]now plug in\[s=-4:A(-4+4)+B(-4)=4\implies -4B=4\implies B=-1\]

OpenStudy (anonymous):

I didnt get those values though :(

OpenStudy (turingtest):

do you see how I got mine?

OpenStudy (anonymous):

i mean for the values of the other one

OpenStudy (turingtest):

that I haven't checked yet, one sec...

OpenStudy (turingtest):

\[A(s+1)(s+4)+Bs(s+4)+Cs(s+1)=2\]\[s=-1:-3B=2\implies B=-\frac23\]\[s=-4:12C=2\implies C=\frac16\]\[s=0:4A=2\implies A=\frac12\]yes I guess you are right :)

OpenStudy (anonymous):

yay!! so would the final answer be (1/2)-(2/3)(e^-t)+(1/6)e^(-4t)-e^(-4t)

OpenStudy (turingtest):

looks good to me :)

OpenStudy (anonymous):

it's wrong :(

OpenStudy (turingtest):

I think you forgot the one from the other term do you have infinite tries?

OpenStudy (anonymous):

no I ownly have 4 more tries

OpenStudy (anonymous):

but what did I forget

OpenStudy (anonymous):

only*

OpenStudy (turingtest):

ok let me try it again from the top; I went of what you had I should double-check everything

OpenStudy (anonymous):

ok thanks

OpenStudy (turingtest):

ok I think I found it...

OpenStudy (turingtest):

\[x''(0)+4x'+8x=2e^{-t}\]\[s^2X(s)-sx(0)-x'(0)+4sX(s)-4x(0)+8X(s)=\frac2{s+1}\]\[(s^2+4s+8)X(s)=\frac2{s+1}+4\]

OpenStudy (turingtest):

you dropped the 8...

OpenStudy (turingtest):

unfortunately this is going to make partial fractions a real pain...

OpenStudy (turingtest):

write is as a complete square\[((s+2)^2+4)X(s)=\frac2{s+1}+4\]you should wind up with some hyperbolic I think

OpenStudy (turingtest):

hyperbolic trig functions*

OpenStudy (anonymous):

can you please show me because I honestly have no clue

OpenStudy (turingtest):

\[X(s)={2\over{((s+2)^2+4)(s+1)}}+\frac4{((s+2)^2+4)}\]\[={2\over{((s+2)^2+4)(s+1)}}+2\frac2{((s+2)^2+2^2)}\]the second term is form 19 on the chart in your link

OpenStudy (turingtest):

\[\mathcal L^{-1}\{\frac2{((s+2)^2+2^2)}\}=e^{-2t}\sin(2t)\]

OpenStudy (turingtest):

the other part... you're gonna have to use partial fractions fractions and it looks like it's gonna be a pain

OpenStudy (turingtest):

\[{2\over{((s+2)^2+4)(s+1)}}={As+B\over(s+2)^2+4}+{C\over s+1}\]\[(As+B)(s+1)+C[(s+2)^2+4]=2\]\[s=-1:5C=2\implies C=\frac25\]\[s=0:B+\frac{16}5=2\implies B=-\frac65\]

OpenStudy (turingtest):

I think you'll have to get A on your own ;) I'm a bit busy

OpenStudy (turingtest):

I think\[(A+C)s^2=0\implies A=-C=-\frac25\]but you shouod double-check me

OpenStudy (turingtest):

I think that is correct